Bond A Maturity Coupon Yield 2 years 6% 6% Price 100

Coupon is paid semi-annually and your investment horizon is 1.5 years. A) What is the modified duration of this bond? B) What is the convexity of this bond? C) If the yield after 6 months is expected to fall to 5.5% and then to 4.5% from the beginning of year 2, what would be your holding period return, price appreciation/depreciation, re-investment income and coupon income from this bond?
Answer #1

Solution

Modified Duration = D / (1+r)

where D= Duration

r = Interest Rate

For Calculation of Duration is the aggregate of weighted cash flows divided by Aggregate weights.

For Bond A , Let us calculate duration for 2 years, semi annually, as in semi annually it will be for 4 periods and the coupon rate of 6% p. a shall be 3 % semi annually, the bonds will redeemed at par.

Periods (a) Cash Flows @ 3 % (b) Discounting Factor @ 6 % (c) Discounted Value (d)= (b)* (c) Weighted Discount Value (e)= (a)*(d) (t2+t) (f)= (a2+a ) g= (f)*(c)
1 3 0.9433 2.83 2.83 2 5.66
2 3 0.8899 2.67 5.34 6 16.02
3 3 0.8396 2.52 7.56 12 30.24
4 3 0.7920 2.38 9.52 20 47.6
4 100 0.7920 79.20 316.8 20 1584
\sumDiscounted Value 89.6 \sumWeighted Discounted Values 342.05 1683.52

Duration = \sum Weighted Discount/ \sum Discounted Value = 342.05/89.6 = 3.84 period , that is if we convert this in yearly , it means 3.84/2 = 1.92 years

Now Modified Duration is = Duration / (1+r) = 1.92/ (1+0.06) = 1.92/1.06 = 1.81 years.

Modified Duration is 1.81 Years

B) Convexity of Bond

Convexity = [1/ P * (1+ y)2]* \sum t=4[ CFt/(1+y)t(t2+t)

where , P = Current Market Price of Bond , which is 89.6

y= YTM of the Bond; 6%

t= no of periods , 4

Now, using the above table's data in the above formula ,

Convexity= [1/ 89.6(1+0.06)2](1683.52) = [1/89.6*1.1236]* 1683.52= [1/100.67]*1683.52= 16.72 years2

Convexity = 16.72 years2

C) If YTM falls by 5.5% after 6 months then current price of the bond will increase and if during the beginning of 2nd year YTM falls further 4.5% then current price will further fall.

If YTM changes to 5.5%, the duration will change , we will use the duration of the period that is 3.84

Change in Price = - D * Change in YTM

= -3.84 * (0.06-0.055)

=-3.84 * 0.005= -0.0192 or -1.92%

New Price = 89.6* (1-0.0192)= 89.6*0.9808= 87.88

If YTM changes to 4.5%, the duration will change , we will use the duration of the period that is 3.84

Change in Price = - D * Change in YTM

= -3.84 * (0.06-0.045)

=-3.84 * 0.015= -0.0576 or -5.76%

New Price = 89.6* (1-0.0576)= 89.6*0.9424= 84.44

Price at the end of 2 nd year = 84.44

Price at the beginning = 89.60

  • Holding Period Return =[( Price at the end - Price at the beggining )/ Price at the beginning ]* 100 = [(84.44-89.60)/89.60]*100= -5.76%
  • Price depreciation of 5.16, as the closing price is less than the current price.
  • Reinvestment Income shall be if interest rate decline to 4.5 % at the end of year 1 , the
Period Cash Flows Reinvestment Rate Period Amount at period end
1 6 6% 4 7.57
2 5.5 5.5% 3 6.42
3 4.5 4.5% 2 4.92
4 4.5 4.5% 1 4.70
4 100 100
123.61
  • Coupon rate shall increase .